Let x denote the time taken to run a road race. Suppose x is approximately normally distributed with a mean of 190 minutes and a standard deviation of 21 minutes. If one runner is selected at random, what is the probability that this runner will complete this road race in 218 to 242 minutes?

Round answer to 4 decimal places.
Answer #1

Solution :

Given that ,

mean = = 190

standard deviation = = 21

P(218 < x < 242) = P[(218 - 190)/ 21) < (x - ) /  < (242 - 190) / 21) ]

= P(1.33 < z < 2.48)

= P(z < 2.48) - P(z < 1.33)

= 0.9934 - 0.9082

= 0.0852

Probability = 0.0852